How can $3,960 be invested, part of it at 4% and the remainder at 5%, so that the two investments will produce the same amount of interest?
Answer by stanbon(75887)

How can $3,960 be invested, part of it at 4% and the remainder at 5%, so that the two investments will produce the same amount of interest?
-------------------
Equation:
0.04x - 0.05(3960-x) = 0
----
4x - 5(3960-x)
---
4x -5*3960 + 5x = 0
9x = 5*3960
x = $2200 (amt. to invest at 4%)
-----------
3960 - x = 3960-2200 = $1760 (amt. to invest at 5%)
============
Cheers,
Stan H.
============